# Finding Counterfactual Evidences for Node Classification

### v1

@@ Abstract.
Counterfactual learning is emerging as an important paradigm, rooted in causality, which promises to alleviate common issues of graph neural networks (GNNs), such as fairness and interpretability. However, as in many real-world application domains where conducting randomized controlled trials is impractical, one has to rely on available observational (factual) data to detect counterfactuals.
In this paper, we introduce and tackle the problem of searching for counterfactual evidences for the GNN-based node classification task. A counterfactual evidence is a pair of nodes such that, regardless they exhibit great similarity both in the features and in their neighborhood subgraph structures, they are classified differently by the GNN. We develop effective and efficient search algorithms and a novel indexing solution that leverages both node features and structural information to identify counterfactual evidences, and generalizes beyond any specific GNN. Through various downstream applications, we demonstrate the potential of counterfactual evidences to enhance fairness and accuracy of GNNs.
copyright: nonedoi: isbn: conference: ; August 3 - 7; Toronto, Canadajournalyear: 2025
@@ 1. Introduction
Besides the general importance of finding counterfactuals in the observational data (previously discussed), there are some immediate applications of graph counterfactual evidence that we develop in case studies in §7. Intuitively, finding two data points which are very similar but treated differently, might represent a violation of the principle that “similar individual should be treated similarly”(Dwork et al., 2012), thus indicating potential unfairness issues with the model. Two similar yet classified differently nodes might also be an indication of an area around the decision boundary of the GNN, where misclassification errors concentrate, thus providing a signal for potential intervention to improve the performance of a classifier.
Our contributions. We tackle the problem of extracting, from a given graph, a pair of nodes (together with their neighborhoods), such that they exhibit high similarity in both structure and features, yet they are classified differently by a GNN. More specifically, we consider two versions of the problem: the local version seeks counterfactual evidences for a given target node, while the global version seeks for counterfactual evidences among all possible node pairs. For both problems, we aim to extract the top-$k$ counterfactual evidences w.r.t. a measure of similarity of the two neighborhood subgraphs. As a measure of similarity, we adapt the Weisfeiler Lehman (WL) kernel-based graph similarity (Shervashidze and Borgwardt, 2009; Shervashidze et al., 2011), for the case of node-anchored graphs (i.e., $L$-hop neighbor subgraphs), having multiple node features. Recall that the WL-test is a necessary but insufficient condition for graph isomorphism (Shervashidze and Borgwardt, 2009), whereas the recent GNN variants are not more powerful than the 1-WL test (Xu et al., 2019). Our technique inherits the simplicity and computational efficiency of the WL kernel computation, while following the update scheme of message-passing GNNs, thus it generalizes the inference process without being tailored for any specific GNN.
@@ 3. Problem statement
Graph Neural Networks (GNNs) (Kipf and Welling, 2017) comprise a well-established family of deep learning models tailored for analyzing graph-structured data. GNNs generally employ a multi-layer message-passing scheme as shown in Equation 1.
@@ 5.1. Local CE Identification
Time and Space Complexity. The time cost of LocalCE-I consists of four parts, with the first three being similar to those in LocalCE-B: (1) Inference cost of GNN: $O\left(Ld\left(|E|+|V_{c}||\mathbf{Y}|\right)\right)$, where $|V_{c}|=\frac{1}{m}|V_{test}|$ indicates the average number of test nodes per cluster; (2) vector aggregation cost for KS score: $O(Ld|V_{c}||N(\cdot)|)$, where $|N(\cdot)|$ is the maximum number of 1-hop neighbors of a test node; (3) cost of finding the top-$k$ results: $O(k|V_{c}|+p)$; and (4) offline index construction cost: $O(mp|V_{test}|(d+{\sf{HS}}^{d}))$, where ${\sf{HS}}^{d}$ is the time for computing the intersection in $d$-dimensional hyperspherical space. Therefore, the online time complexity of LocalCE-I is roughly reduced by a factor of the number of clusters.
For space complexity, the additional index overhead is $O(mp|V_{test}|)$.
@@ 5.2. Global CE Identification
The identification of global CEs is a natural extension of the local algorithms. As shown in Algorithm 3 (in Appendix), we select one local algorithm (LocalCE-B or LocalCE-I) to retrieve the top-$1$ LCE for each test node, and then select the best one as the top-$1$ GCE. As for the extension to top-$k$ GCEs, we maintain a bucket $B$ of size $k$, and incrementally add the top-$1$ LCE among all test nodes.
@@ 6. Experimental Results
We conduct experiments to demonstrate the effectiveness, efficiency, scalability, and generalizability of our solutions for finding both local and global counterfactual evidences. Our algorithms are implemented in Python 3.10.14 by PyTorch-Geometric framework. All experiments are conducted on the single core of a Linux system equipped with AMD EPYC 7302P CPU and 256 GB RAM. Our code and data are available at (cod, 2024).
@@ 6.1. Experimental Setup
Datasets. We utilize datasets from various real-world domains to showcase the performance of our methods. The statistics of the datasets are shown in Table 1. The details are in the Appendix (§ A).
@@ 7.2. Verifying Prediction Errors
We employ the classic GCN (Kipf and Welling, 2017) as the classifier and use two widely used graph node classification datasets: Cora (McCallum et al., 2000) and PubMed (Sen et al., 2008). We evaluate the effectiveness of CEs based on the accuracy of the classifier, where the accuracy is the ratio of the number of correctly classified nodes to the total number of nodes.
@@ 8. Conclusions
Appendix A Dataset details
Appendix B Additional Case Studies

### v2

@@ Abstract.
Counterfactual learning is emerging as an important paradigm, rooted in causality, which promises to alleviate common issues of graph neural networks (GNNs), such as fairness and interpretability. However, as in many real-world application domains where conducting randomized controlled trials is impractical, one has to rely on available observational (factual) data to detect counterfactuals. In this paper, we introduce and tackle the problem of searching for counterfactual evidences for the GNN-based node classification task. A counterfactual evidence is a pair of nodes such that, regardless they exhibit great similarity both in the features and in their neighborhood subgraph structures, they are classified differently by the GNN. We develop effective and efficient search algorithms and a novel indexing solution that leverages both node features and structural information to identify counterfactual evidences, and generalizes beyond any specific GNN. Through various downstream applications, we demonstrate the potential of counterfactual evidences to enhance fairness and accuracy of GNNs.
counterfactual evidence, node classification, graph neural networks
conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736960isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Graph-based database modelsccs: Information systems Database query processing
@@ 1. Introduction
Besides the general importance of finding counterfactuals in the observational data (previously discussed), there are some immediate applications of graph counterfactual evidence that we develop in case studies in §7. From the standpoint of a watchdog auditing an algorithmic decision-making system for unlawful discrimination, it is critical to find evidences of nodes which are similar, but treated differently, thus violating the principle that “similar individuals should be treated similarly” (Dwork et al., 2012), indicating potential unfairness issues with the model. Two similar yet classified differently nodes might also be an indication of an area around the decision boundary of the GNN, where misclassification errors concentrate, thus providing a signal for potential intervention to improve the performance of a classifier.
Our contributions. We tackle the problem of extracting, from a given graph, a pair of nodes (together with their neighborhoods), such that they exhibit high similarity in both structure and features, yet they are classified differently by a GNN. Our is a data mining problem, that is, we define a novel structure of interest in a large graph based on a pre-trained GNN model’s output and we then devise the algorithm to extract it. More specifically, we consider two versions of the problem: the local version seeks counterfactual evidences for a given target node, while the global version seeks for counterfactual evidences among all possible node pairs. For both problems, we aim to extract the top-$k$ counterfactual evidences w.r.t. a measure of similarity of the two neighborhood subgraphs. As a measure of similarity, we adapt the Weisfeiler Lehman (WL) kernel-based graph similarity (Shervashidze and Borgwardt, 2009; Shervashidze et al., 2011), for the case of node-anchored graphs (i.e., $L$-hop neighbor subgraphs), having multiple node features. Recall that the WL-test is a necessary but insufficient condition for graph isomorphism (Shervashidze and Borgwardt, 2009), whereas the recent GNN variants are not more powerful than the 1-WL test (Xu et al., 2019). Our technique inherits the simplicity and computational efficiency of the WL kernel computation, while following the update scheme of message-passing GNNs, thus it generalizes the inference process without being tailored for any specific GNN.
@@ 3. Problem statement
Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Du et al., 2021; Qiu et al., 2025) comprise a well-established family of deep learning models tailored for analyzing graph-structured data. GNNs generally employ a multi-layer message-passing scheme as shown in Equation 1.
@@ 5.1. Local CE Identification
Time and Space Complexity. The time cost of LocalCE-I consists of four parts, with the first three being similar to those in LocalCE-B: (1) Inference cost of GNN: $O\left(Ld\left(|E|+|V_{c}||\mathbf{Y}|\right)\right)$, where $|V_{c}|=\frac{1}{m}|V_{test}|$ indicates the average number of test nodes per cluster; (2) vector aggregation cost for KS score: $O(Ld|V_{c}||N(\cdot)|)$, where $|N(\cdot)|$ is the maximum number of 1-hop neighbors of a test node; (3) cost of finding the top-$k$ results: $O(k|V_{c}|+p)$; and (4) offline index construction cost: $O(mp|V_{test}|(d+{\sf{HS}}^{d}))$, where ${\sf{HS}}^{d}$ is the time for computing the intersection in $d$-dimensional hyperspherical space (Lee and Kim, 2014). Therefore, the online time complexity of LocalCE-I is roughly reduced by a factor of the number of clusters.
For space complexity, the additional index overhead is $O(p|V_{test}|)$.
Algorithm 3 GlobalCE-B&I
0:  Graph $G$, GNN $M$, test set $V_{test}$, set of partitions $P$.
0:  Top-$1$ global counterfactual evidence ${\sf GCE}_{opt}(V_{test})$.
1:  Select LocalCE-B or LocalCE-I for local CE identification.
2:  Identify top-$1$ LCE for each test node based on the selected local algorithm.
3:  Identify top-$1$ GCE among the top-$1$ LCEs.
4:  return  ${\sf GCE}_{opt}(V_{test})$.
@@ 5.2. Global CE Identification
The identification of global CEs is a natural extension of the local algorithms. As shown in Algorithm 3 , we select one local algorithm (LocalCE-B or LocalCE-I) to retrieve the top-$1$ LCE for each test node, and then select the best one as the top-$1$ GCE. As for the extension to top-$k$ GCEs, we maintain a bucket $B$ of size $k$, and incrementally add the top-$1$ LCE among all test nodes.
@@ 6. Experimental Results
We conduct experiments to demonstrate the effectiveness, efficiency, scalability, and generalizability of our solutions for finding both local and global counterfactual evidences. Our algorithms are implemented in Python 3.10.14 by PyTorch-Geometric framework. All experiments are conducted on the single core of a Linux system equipped with AMD EPYC 7302P CPU and 256 GB RAM. Our code and data are available at (cod, 2025).
@@ 6.1. Experimental Setup
Datasets. We utilize datasets from various real-world domains to showcase the performance of our methods. The statistics of the datasets are shown in Table 1.
@@ 7.2. Verifying Prediction Errors
We employ the classic GCN as the classifier and use two widely used graph node classification datasets: Cora (McCallum et al., 2000) and PubMed (Sen et al., 2008). We evaluate the effectiveness of CEs based on the accuracy of the classifier, where the accuracy is the ratio of the number of correctly classified nodes to the total number of nodes.
@@ 8. Conclusions
Acknowledgements.
Dazhuo Qiu and Arijit Khan acknowledge support from the Novo Nordisk Foundation grant NNF22OC0072415. Arijit Khan and Yan Zhao are corresponding authors.
Appendix A Additional Case Studies